OpenStudy (anonymous):

Does the amount of molecules of gas change at a temp lower than STP? the temperature is 175 K.

OpenStudy (anonymous):

Urgent have Finals tomorrow.

OpenStudy (photon336):

Well, @Clemsongator at lower temperatures gas molecules would be moving around much more slowly.

OpenStudy (photon336):

but in terms of the amount of gas, generally it would be constant.

OpenStudy (anonymous):

so the amount would be about the same as STP?

OpenStudy (photon336):

Yeah, remember you said lower temperature. The lower the temperature, the slower the gas molecules move. remember temperature = the average kinetic energy, of the particles. 0.5mv^2. let's look at the ideal gas equation. pV = nRT usually the number of moles of gas n and R are usually constants.

OpenStudy (anonymous):

okay so if have 24.23L of nitrogen monoxide at STP I would still have that amount?

OpenStudy (anonymous):

okay

OpenStudy (photon336):

\[\frac{ 1 mol of any gas }{ STP } = 22.4L \]

OpenStudy (photon336):

if we change the temperature, the volume that the gas will occupy will be different, but the number of moles will be the same.

OpenStudy (anonymous):

okay the question wants to know how many molecules would be in the sample at 175K

OpenStudy (photon336):

okay let me see

OpenStudy (photon336):

okay so if have 24.23L of nitrogen monoxide at STP. So sw know at STP 1 mol of any gas occupies 22.4L right? so our gas occupies 24.32 liters so it must be a little bit over 1 mole. see below. we setup a porportion. \[\frac{ mol }{ 22.4L } = \frac{ x }{ 24.32L } = \frac{ 24.32 }{ 22.4 } = 1.1 mol, of NO\] okay the question wants to know how many molecules would be in the sample at 175K, This is our inforamtion that we have. 273K, 1atm, 1.1 mol, 24.32L when we lower the temperature, we may cause a change in the pressure or volume, but not the number of moles. to do calcualtions with that equation we must keep one of the variables constant. pV = nRT essentially the number of moles should be the same. essentially this equation rests on the fact that we study changes in pressure temperatuer and volume, with the idea that the number of moles is constnat.

OpenStudy (anonymous):

Thanks

OpenStudy (photon336):

@clemsongator np